London Switch Updates
Resolved Dec 04 at 05:00pm GMT
We will be performing maintenance to one of the switches in our client cabinets.
During the window we intend to failover customers to our relay switch in that space. There will be a brief interruption to services as we switch over.
Resolved Nov 30 at 02:00pm GMT
We will be performing maintenance on our Portal API to better prepare for our Cloud Instance product release.
During this time there may be a period where tickets and billing with be unavailable. All client services will be unaffected.
Resolved Nov 01 at 08:52am GMT
While debugging an issue with recaptcha on our portal, the API was restarted several times resulting in some unavailability.
We have found the root cause of the problem to be with DNS resolving of the google.com domain and have implemented a fix. We do not expect any further downtime.
IPv4 Traffic Dropped
Resolved Oct 30 at 09:38am GMT
08:53AM we were notified of an outage of all client services in our London location.
08:55AM as30456 noc team were notified of the outage requesting assistance.
09:02AM we identified the cause of the issue being all ipv4 prefixes within London had been dropped.
09:11AM we received acknowledgement from an engineer at as30456 investigating the issue.
09:17AM our ipv4 prefixes were re-routed after a soft bgp session refresh on as30456-side
09:18AM we received notificatio...
3 previous updates
London service move (on-site)
Resolved Oct 28 at 10:30am BST
On Saturday 28th October we will be performing a move to select customers based in our London location.
This is to allow us to better utilize existing space and power restrictions in our existing space. Affected clients will be emailed to notify them that they will be affected by the end of day on Tuesday 24th October.
Any individual service will not be down for longer than 5-10 minutes at any given time. Services will be moved one-by-one and their connectivity status confirmed after rec...
Resolved Oct 26 at 04:24am BST
Performing some routine API maintenance
Resolved Oct 01 at 04:25pm BST
Maintenance has been complete.
As part of the new firewall overhaul, rules have not been migrated.
If you wish to have your rules migrated, please make a ticket.
1 previous update